OpenStudy (keana):

solve 6/x+4=1/2 a: 8 b: -1 c: 16 d: 4

OpenStudy (australopithecus):

do you mean? \[\frac{6}{x} + 4 = \frac{1}{2}\]

hero (hero):

Hint: Cross Multiply to get 6(2) = x + 4

OpenStudy (keana):

@Australopithecus nope

OpenStudy (australopithecus):

\[\frac{6}{x} = \frac{1}{2} - 4\] \[(\frac{6}{x})x = (\frac{1}{2} - 4)x\] \[6 = (\frac{1}{2} - 4)x\] \[\frac{x}{(\frac{1}{2}-4)} = \frac{(\frac{1}{2} - 4)x}{(\frac{1}{2}-4)}\]

OpenStudy (australopithecus):

oops

OpenStudy (australopithecus):

\[\frac{6}{\frac{1}{2} - 4} = x\]

OpenStudy (australopithecus):

that last equation should have a 6 above it instead of an x

OpenStudy (keana):

@Australopithecus i'm confused

hero (hero):

@Australopithecus made an incorrect assumption it seems

hero (hero):

\(\huge\frac{6}{x+4} = \frac{1}{2}\)

hero (hero):

^Cross multiply that to get 6(2) = x + 4 Then solve for x

OpenStudy (keana):

@Hero so A ?

hero (hero):

Correct @keana

OpenStudy (australopithecus):

\[\frac{6}{x+4} = \frac{1}{2}\] \[\frac{6}{x+4}(x+4) = \frac{1}{2}(x+4)\] \[\frac{6}{x+4}\frac{(x+4)}{1} = \frac{1}{2}\frac{(x+4)}{1}\] \[\frac{6(x+4)}{1*(x+4)} = \frac{(x+4)*1}{2*1}\] \[6 = \frac{x+4}{2}\] \[6(2) = \frac{(x+4)2}{2*1}\] \[6(2) = x+4\] \[6(2) - 4 = x\]